• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name E. Schulte

    Log entry time 15:55:30 on November 08, 2008

    Entry number 247609

    keyword=Day Shift Summary, November 8, 2008

    08:00 Start shift taking production data on 3He target, beam current 8uA,
    HRS-L P=2.35, theta=16.003, BigBite Sieve OUT, Target Collimator IN! ROC
    15 died. It was rebooted. Must get from Alexandre when he trundles past
    how to reboot it.

    09:14 MCC called to let us know that they need to take the beam away to
    work on a cooling water leak.

    09:17 Controlled Access during the water leak fix. Alexandre went down
    to look at the RICH to try to fix the dead time.

    10:14 Checked with MCC about the accelerator status. They are out of the
    accelerator and are starting to get their magnets turned back on.
    Alexandre is leaving the hall, and we will be going back to Beam Permit
    once he is out.

    10:23 Moving target to empty to check beam positions on return to beam
    delivery. And then we decided to move to the BeO target so that MCC
    could see the beam.

    10:44 MCC called to let us know that they have a magnet that is not
    recovering properly from the water leak fix. They estimate about half an
    hour, but will call us again as they know more. The expert is on hi or
    her way.

    11:24 MCC called to start sending pulsed beam.

    11:27 MCC is going to mask beam and put in the 3He target.

    11:29 Requested beam back with raster ON!! from MCC. Requested 2uA to
    check everything out.

    11:37 Spot++ looked good and we requested 8uA. MCC is working on getting
    the FFB working.

    11:40 Back to production running with FFB on and the requested 8uA of
    current.

    12:03 MCC caleld to let us know that they've been having RF issues that
    they're having difficulty recovering from. This is causing long down
    times when they have the problem.

    13:23 We've been running along here. The beam's been pretty good, a
    little trippy, but not too bad. Of course, as I make note of this, the
    beam has tripped and been off for a minute or so...

    14:36 The BTA is being particularly strange at recording the beam used
    today. We've been taking data straight through since return of beam, but
    it only gives us credit for about 1/3 of it. Very odd.

    15:10 A shift in the timing spectrum was noticed for run 3063. Kalyan is
    going to ask Vince about it.

    15:55 Shift ends taking data reasonably smoothly!